OpenStudy (anonymous):

how to find the domain for xsquareroot(4-x^2)?

OpenStudy (amistre64):

well, 4-x^2 cant be less than zero.

OpenStudy (amistre64):

sqrt(-#) is verbotten

OpenStudy (amistre64):

4-x^2 >= 0 -x^2 >= -4 x^2 <= 4 x <= +- sqrt(4) x <= -2 and x <=2 but its wise to double check ;)

OpenStudy (amistre64):

Dom: [-2,2] looks about right to me

OpenStudy (amistre64):

<........-2............2............> - 0 + 0 - are our results right?

OpenStudy (anonymous):

\[-2lex \ge\]

OpenStudy (anonymous):

oops dw about that i pressed s9mething wrong

OpenStudy (anonymous):

THANK YOU AMISTRE64